# Non-Unitary Quantum Many-Body Dynamics using the Faber Polynomial Method

### v2

@@ 4.2 Domain Wall Melting for Hatano-Nelson
We can provide a simple argument to justify the renormalisation of the velocity due to non-hermiticity. The renormalised velocity can be obtained by using a local continuity equation for the particle density (Eq. 23) by working the term $\mathcal{T}_{\ell}$. In the non-interacting limit, this term can be expanded using Wick’s theorem in a local and non-local term,
[…]
The second term of this equation is highly non-local, and it should be zero outside the light-cone region. We proceed by writing the local continuity equation for a site $\ell$, outside the light-cone region, for an instant of time before the arrival of the magnetisation wavefront. Under these conditions $I_{\ell}=0$ and $\langle c^{\dagger}_{\ell}c_{\ell}=0\rangle$ for $\ell>0$ or $I_{\ell-1}=0$ and $\langle c^{\dagger}_{\ell}c_{\ell}=1\rangle$ for $\ell<0$. Furthermore, the non-local correlator vanishes, and so the continuity equation is approximately given by
@@
The $\pm$ sign reflects the different directions of propagation. This formula holds only before the particle density wavefront reaches this site. As after, there is the development of off-diagonal correlations within the light-cone area.
Since at short times we can still identify a sharp light cone, it is tempting to try to formulate a hydrodynamic description for the HN model. Although some advances have been made within the framework of Linblad dynamics  [95, 96], formulating a hydrodynamic description for the non-Hermitian variant of the XXZ model remains extremely challenging, even in the absence of interactions. The nonlinearity of the equations of motion results in the loss of most local conservation laws. Surprisingly, as we show for the Hatano-Nelson model in Fig. 5, the hydrodynamic equations still hold in the initial time. To make a concrete comparison between the numerics and the analytical predictions, we proceed in a somewhat phenomenological way and incorporate the renormalised velocity into the hydrodynamic expressions derived for the Hermitian case [97].
@@
$I\left(\ell,t\right)=\dfrac{1}{\pi}\sqrt{1-\left(\dfrac{\ell}{v_{\text{eff}}\;t}\right)^{2}}.$ (29)
@@
This length is marked by the white dashed lines in the Fig. 4. We just stress that this reasoning is only valid for $\gamma>0$, since for $\gamma<0$ the magnetisation wavefront necessarily reaches the system boundary. The expression suggests that faster quasiparticles with velocity $v^{\ast}=J+\gamma$ cease the ballistic propagation of the magnetisation wavefront upon hitting the physical boundary. Thus, the maximum propagation velocity allowed in the system increases compared to the Hermitian case, since particles can hop from right to left with velocity $J+\gamma$ (assuming $\gamma>0$). Following the initial propagation of the magnetisation wave front, the system reaches a steady state characterised by the presence of a charge current traversing the two domains. This current emerges due to a flux of particles sourced from the environment, which are subsequently annihilated, as depicted in Fig. 5. However, contrary to the setting described by Kawabata [34], non-Hermiticity leads to the spatial suppression of this current.
@@ 5 Application to Non-Hermitian Many-Body Systems
Figure 8: Long-time behaviour of the magnetisation profile shown in Fig. 7, $\langle S_{\ell}^{z}\rangle_{\infty}$, for different values of $\Delta$. We see that at $\Delta=0$ an emergent domain wall is formed, which is stable for small $\Delta$ (left panel). Increasing $\Delta$ the system develops a potential drop in the middle of the chain, akin to diffusive dynamics, which then further develops an oscillating patter as $\Delta$ increases (right panel). Other parameters $\gamma=0.8J$.
@@ 5.1 Magnetisation Dynamics in the Interacting Hatano-Nelson Model
where $\tau^{\ast}$ is the initial time corresponding to the reshuffling of magnetic until there is a stable boundary accumulation. We plot in Fig. 8 $\left\langle S^{z}_{\ell}\right\rangle_{\infty}$ for different values of $\Delta$. We see that the emergent domain-wall state generated by the non-reciprocal exchange is stable at weak interactions (left panel). However, as $\Delta$ increases, a novel region emerges that interpolates between the two magnetic domains (right panel). In particular, we see that the system develops a magnetisation drop, similar to a potential drop in systems that show diffusive transport. Upon further increasing $\Delta$ we see the emergence of antiferromagnetic correlations on top of this magnetisation slope, which, as expected, are frozen by the large interaction and do not decay away. The result above is particularly intriguing as it suggests that non-reciprocal coupling and interaction collaborate to establish a current-carrying steady state in the centre of the system: the former driving the formation of a domain wall that acts as source and drain, the latter providing the necessary scattering term to dissipate and establish a finite average current. In Fig. 9 we plot the spatio-temporal dynamics of the local current profile as a function of interaction and non-reciprocity. Comparable findings have been reported [102] for the ground state of the interacting Hatano-Nelson model with nearest-neighbour repulsion. In that study, they also noted that the initial magnetisation profile (referred to as the real-space Fermi surface) is disrupted by interactions that induce a Néel order. In this work, we demonstrate that such a phenomenon can also be dynamically generated by the non-unitary time-evolution.
Similarly to the non-interacting case, there is a current in this interpolating region, which satisfies the same continuity equation as in Eq. 27. As clearly seen in Fig. 9, once again there is a competition between non-reciprocity and the interaction parameter $\Delta$ in defining the size of this intermediate region, where it is possible for the particle to enter from the environment and give rise to this current. As one is looking at small system sizes, it is not possible to create a stable steady current (like in the non-interacting case). Due to this, we see oscillations in the direction of the central current, and thus there is current flowing in both directions. However, in the extreme case where $\gamma=J$, this oscillatory behaviour ceases to exist, and the current has a fixed direction. When $\Delta$ is much greater than $J$, the current disappears in the bulk region, where the Néel order is maintained.
@@ 5.2 Effect of interaction in the Domain-Wall melting for Hatano-Nelson
Figure 11: Magnetisation profile in function of the lattice site and time. Left - Dynamics with the jump operator, $\sqrt{\left|\gamma\right|}\left(S^{-}_{\ell}-i\text{sgn}\left(\gamma\right)S^{-}_{\ell+1}\right)$. Right - Dynamics with the jump operators, $\sqrt{\gamma}S^{+}_{\ell}S^{-}_{\ell+1}$. The system was initially prepared in a Néel State (Eq. 33). Other parameters: $\Delta=0$, $\gamma=0.8J$ and $L=20$.
@@ 6 Application to Quantum Jumps Unravelling
The quantum jump is applied by first selecting the quantum jump channel in accordance to their probability,
@@
This algorithm gives access to the full Lindbladian dynamics, by averaging over the propagated quantum trajectories. Further, it also provides access to the dynamics under continuous monitoring [20, 9]. This is achieved by following the many-body quantum trajectory and computing, for example, non-linear functions of the state. One such function is the entanglement entropy of quantum trajectories, as we shall discuss below.
@@ 6.1 Magnetisation and Entanglement Dynamics in Monitored Spin Chains
where $J$ is the XX exchange term between neighbouring spins and $\Delta$ an Ising like exchange. We compare the dissipative dynamics generated by two different types of jump operators. A first set of jumps that we consider describe next-neighbour decoherence of spin excitations along the chain and take the form
@@
$L_{1+\ell}$ $=\sqrt{\left|\gamma\right|}\left(S^{-}_{\ell}-i\text{sgn}\left(\gamma\right)S^{-}_{\ell+1}\right),\;\ell=\{0,\cdots,L-2\}$
@@
Interestingly, with these choice of jump operators the non-Hermitian Hamiltonian associated to the no-click limit turns out to be given by a spin version of the many-body Hatano-Nelson model. Indeed, the following straightforward calculation yields the non-Hermitian Hamiltonian
@@
The last term does not affect the dynamics, as it just an overall background decay. Although this model can be connected to the fermionic version of the Hatano-Nelson model, the Linblad equation is not quadratic due to the Jordan-Wigner strings in the terms of the form $S^{-}\rho S^{+}$.
@@
with =$\ell=\left\{0,\cdots,L-2\right\}$. Previous studies have shown that these operators induce a phenomenon known as the Liouvillian skin effect [24, 105, 106, 107]. Specifically, there is an exponential localisation of the Liouvillian modes at the boundaries of the system. This becomes evident when the Linblad is projected onto the one-particle sector, where, at $J=0$ and $\Delta=0$, it simplifies to an effective Hatano-Nelson Hamiltonian tuned to its exceptional point [105],
@@
$\mathcal{H}_{\text{eff}}&=-\frac{J}{2}\sum_{\ell=0}^{L-2}\left[S_{\ell}^{+}S_{\ell+1}^{-}+\text{h.c}\right]-\left[\Delta-i\dfrac{\gamma}{2}\right]\sum_{\ell=0}^{L-2}S_{\ell}^{z}S_{\ell+1}^{z}\\ &\quad+\frac{i\gamma}{4}\left(S_{0}^{z}-S_{L-1}^{z}\right)-\frac{i\gamma}{8}\left(L-1\right)\end{aligned}.$ (44)
@@
In Fig. 11, we plot the magnetisation dynamics starting from a product state and evolving under the two types of dissipative evolutions. In the left panel, we plot the dynamics generated by Eq. 40. Under this set of jump operators, manifestations of non-reciprocity and charge accumulation are still present in the transient dynamics [108, 109]. We note that these are the same jump operators considered in [109] with the phase, $\phi$, defined by them, tuned to get the non-reciprocal regime. However, in the long time limit, a generic state converges to a configuration with zero excitations, $\ket{\downarrow\cdots\downarrow}$ [109] (see the left plot of Fig. 11). This is attributed to the recycling terms of the form $c_{n}\rho c^{\dagger}_{n}$, which effectively remove particles from the system. Additionally, the system has another dark state within the one-magnon sector, which only plays a role in decelerating the relaxation dynamics towards the fully polarised down state [109]. The final state obtained under this Linbladian does not resemble at all the magnetisation profile that one would have obtained in the no-click limit.
@@
We then focus on the jump operators in Eq. 42 and plot in Fig. 11 (right panel) the resulting magnetisation dynamics. For the sake of simplicity, we consider the dynamics with $\Delta=0$. The dynamics driven by this model will facilitate the accumulation of spins in an up state at the left extremity of the chain. The imaginary Ising exchange term is responsible for diminishing the state’s norm when adjacent spins are antialigned, precipitating a quantum jump that propagates a spin towards the left edge. This phenomenon is depicted in the spatial and temporal evolution of the magnetisation profile shown in Fig. 11. In contrast to the Hatano-Nelson model, this model does not converge to a steady state characterised by zero excitations. This distinction is attributable to a difference in symmetry. The Hatano-Nelson model exhibits only weak U(1) symmetry, whereas the present model possesses strong U(1) symmetry, thereby ensuring that the system state is kept within the initial magnetisation sector at all times.
@@
Figure 12: Top - Time dependence of the conditional entanglement entropy for different system sizes with $\gamma=0.8J$. On the left, the dynamics corresponds to the jump operator $\sqrt{\left|\gamma\right|}\left(S^{-}_{\ell}-i\text{sgn}\left(\gamma\right)S^{-}_{\ell+1}\right)$, while on the right to $\sqrt{\gamma}S^{+}_{\ell}S^{-}_{\ell+1}$. Bottom Left: Time evolution of the entanglement entropy for $\gamma=0.1J$. The inset corresponds to the fit of the steady-state entanglement entropy to the law $S_{L/2}\left(\infty\right)=a_{0}+a_{1}L$, with $a_{0}=\left(0.145\pm 0.002\right)$ and $a_{1}=\left(0.18\pm 0.03\right)$. Bottom Right - Steady-state entanglement entropy for different system sizes as a function of the non-hermitian parameter. These systems were studied under the jump operators $\sqrt{\gamma}S^{+}_{\ell}S^{-}_{\ell+1}$. Other parameter: $\Delta=0.0J$.
@@
$\bar{S}_{\ell}\left(t\right)=\int\mathcal{D}\xi_{t}\;\mathcal{P}\left(\xi_{t}\right)S_{\ell}\left(\xi_{t}\right).$ (45)
@@
with $S\left(\xi_{t}\right)$ is the entanglement of a given quantum trajectory that evolves according to Eq. 1.
@@
For early times, the entanglement entropy in both models increases linearly. Without non-Hermiticity, the system evolves under standard hermitian dynamics, leading to the entanglement entropy saturating at a value proportional to the system’s volume as it locally thermalises [110, 111]. However, in the presence of non-Hermiticity and quantum jumps, this behaviour may drastically alter. In the Hatano-Nelson model with jumps, the entanglement entropy drops to zero after an initial period determined by the measurement rate $\gamma$. Each jump causes a spin-flip, driving the system to a state devoid of magnetic excitations. This occurs regardless of the system’s total size, resulting in a trivial entanglement area law for any nonzero value of $\gamma$. The no-click limit of the Hatano-Nelson model also has this area law scaling of the entanglement entropy; however, this is driven by the single-particle skin effect[34].
@@
In contrast, the model described by Eq. 44 does not relax to a zero excitation state, as seen in Fig. 12. The strong U(1) symmetry confines the dynamics to the magnetisation sector of the initial state [^3] The dynamics conserves the total number of excitations in the initial state as the initial state is eigenstate of $\sum_{n=0}^{L-1}S^{z}_{n}$ and the Hamiltonian in Eq. 44 as U(1) symmetry.. Initially, the system has a linear growth of the entanglement entropy, reminiscent of the unitary evolution. However, similarly to the Hatano-Nelson model, the steady-state supports an area law entanglement for a positive value of $\gamma$, as seen in the left plot of Fig.12. This can be understood through the spectral properties of the no-click Hamiltonian (Eq. 44). The spectrum is always complex-valued, and so, in the no-click limit, the steady-state corresponds to right-eigenstate with the slowest decaying mode in the zero magnetisation sector. In particular, for certain values of $\gamma/J$, the imaginary component of the spectrum is gapped. Thus, the entanglement entropy inevitably follows an area law in the long-time limit, similar to the ground state of a gapped Hamiltonian [112, 32]. The imaginary gap ($\Delta_{\rm Im}$) in the zero magnetisation sector can be analytically obtained in the limit of $\gamma\gg J$,
$\Delta_{\rm Im}=-i\dfrac{\gamma}{2J}+\mathcal{O}\left(\dfrac{J}{\gamma}\right)$ (46)
@@
The measurement apparatus effectively disentangles the system, pushing excitations towards the left boundary, and preventing the formation of long-range correlations. It is harder to completely confirm this for smaller values of $\gamma$, where the unitary evolution dominates both the non-Hermitian and stochastic terms. This mainly affects systems with a smaller dimension, where finite-size effects are substantial, since the terms proportional to $L^{-n}$ with $n>0$ in the entanglement entropy cannot be ignored. For example, for $\gamma=0.1J$, one cannot extrapolate the true entanglement scaling, as the observed linear growth might be an artefact of small system sizes. Nevertheless, the remaining data points clearly indicate the collapse of all system sizes to the same value, thus revealing the area law nature of the entanglement entropy.

### v3

@@ 4.2 Domain Wall Melting for Hatano-Nelson
A simple argument to justify the renormalisation of the velocity due to non-hermiticity can be provided. The renormalised velocity can be obtained by using a local continuity equation for the particle density (Eq. (23)) and expanding the term $\mathcal{T}_{\ell}$. In the non-interacting limit, this term can be expanded using Wick’s theorem in a local and non-local term,
$\mathcal{T}_{\ell}=$ $-\dfrac{\gamma}{J}\left(I_{\ell}+I_{\ell-1}\right)\left(1-2\left\langle c_{\ell}^{\dagger}c_{\ell}\right\rangle\right)+$ (26)
$-i\gamma\sum_{n\neq\{\ell,\ell-1\}}^{L-2}\left[\left\langle c_{\ell}^{\dagger}c_{n+1}\right\rangle\left\langle c_{n}^{\dagger}c_{\ell}\right\rangle-\left\langle c_{n+1}^{\dagger}c_{\ell}\right\rangle\left\langle c_{\ell}^{\dagger}c_{n}\right\rangle\right].$
The second term of this equation is highly non-local and it is zero outside the light-cone region. We proceed by writing the local continuity equation for a site $\ell$, outside the light-cone region, for an instant of time before the arrival of the particle density wavefront. Under these conditions, $I_{\ell}=0$ and $\langle c^{\dagger}_{\ell}c_{\ell}\rangle=0$ for $\ell>0$; and $I_{\ell-1}=0$ and $\langle c^{\dagger}_{\ell}c_{\ell}\rangle=1$ for $\ell<0$. Furthermore, the non-local correlator vanishes, and so, the continuity equation is approximately given by
@@
The $\pm$ sign reflects the different propagation directions. This formula holds only before the particle density wavefront reaches a given site $\ell$. Afterwards, off-diagonal correlations emerge within the light-cone area.
Since at short times we can still identify a sharp light cone, it is tempting to try to use a hydrodynamic description for the HN model. Although some advances have been made within the framework of Linblad dynamics  [98, 99], formulating a hydrodynamic description for the non-Hermitian variant of the XXZ model remains extremely challenging, even in the absence of interactions. The nonlinearity of the equations of motion results in the loss of most local conservation laws. We proceed in a somewhat phenomenological way and incorporate the renormalised velocity into the hydrodynamic expressions derived for the Hermitian case [100]. Surprisingly, as we show for the Hatano-Nelson model in Fig. 5, the hydrodynamic equations, derived in the unitary problem, still hold in the initial time, as long as the appropriate renormalised velocity is included.
@@
$I\left(\ell,t\right)=\dfrac{1}{\pi}\sqrt{1-\left(\dfrac{\ell}{v_{\text{eff}}\;t}\right)^{2}},$ (29)
@@
It is important to emphasise that these equations only correctly predict the behaviour at short times, where there is ballistic propagation of the particle density wavefront. At later times, the non-hermiticity plays a greater role than merely renormalising the velocity of propagation.
@@
This length is marked by the white dashed lines in Fig. 4. Note that this reasoning is only valid for $\gamma>0$, since for $\gamma<0$ the particle density wavefront necessarily reaches the system’s boundary. The expression suggests that faster quasiparticles with velocity $v^{\ast}=J+\gamma$ cease the ballistic propagation of the particle density wavefront upon hitting the physical boundary. Thus, the maximum propagation velocity allowed in the system increases compared to the Hermitian case, since particles can hop from right to left with velocity $J+\gamma$ (assuming $\gamma>0$). Following the initial propagation of the particle density wavefront, the system reaches a steady state characterised by the presence of a charge current traversing the two domains. This current emerges due to a flux of particles sourced from the environment, which are subsequently annihilated, as depicted in Fig. 5. However, contrary to the setting described by Kawabata [34], non-Hermiticity leads to the spatial suppression of this current.
@@ 5 Application to Non-Hermitian Many-Body Systems
Figure 8: Long-time behaviour of the magnetisation profile shown in Fig. 7, $\langle S_{\ell}^{z}\rangle_{\infty}$, for different values of $\Delta$. We see that at $\Delta=0$ an emergent domain wall is formed, which is stable for small $\Delta$ (left panel). Increasing $\Delta$, the system develops a potential drop in the middle of the chain, akin to diffusive dynamics, which then further develops into an oscillating patter as $\Delta$ increases (right panel). Other parameters $\gamma=0.8J$.
@@ 5.1 Magnetisation Dynamics in the Interacting Hatano-Nelson Model
where $\tau^{\ast}$ is the initial time corresponding to the reshuffling of magnetisation until there is a stable boundary accumulation. We plot in Fig. 8 $\left\langle S^{z}_{\ell}\right\rangle_{\infty}$ for different values of $\Delta$. We see that the emergent domain-wall state generated by the non-reciprocal exchange is stable at weak interactions (left panel). However, as $\Delta$ increases, a novel region emerges that interpolates between the two magnetic domains (right panel). In particular, the system develops a magnetisation drop, similar to a potential drop in systems that show diffusive transport. Further increasing $\Delta$, antiferromagnetic correlations emerge on top of this magnetisation slope, which, as expected, are frozen by the large interaction and do not decay away. The result above is particularly intriguing as it suggests that non-reciprocal coupling and interaction collaborate to establish a current-carrying steady state in the system’s centre: the former driving the formation of a domain wall that acts as source and drain, the latter providing the necessary scattering term to dissipate and establish a finite average current. Comparable findings have been reported [105] for the ground state of the interacting Hatano-Nelson model with nearest-neighbour repulsion. In that study, they also noted that the initial magnetisation profile (referred to as the real-space Fermi surface) is disrupted by interactions that induce a Néel order. In this work, we demonstrate that such a phenomenon can also be dynamically generated by the non-unitary time-evolution.
Similarly to the non-interacting case, there is a current in this interpolating region, which satisfies the same continuity equation as in Eq. (27). As clearly seen in Fig. 9, once again there is a competition between non-reciprocity and the interaction parameter $\Delta$ in defining the size of this intermediate region, where it is possible for the particle to enter from the environment and give rise to this current. As one is looking at small system sizes, it is not possible to create a stable steady current (like in the non-interacting case). Due to this, we see oscillations in the direction of the central current, and thus there is current flowing in both directions. However, in the extreme case where $\gamma=J$, this oscillatory behaviour ceases to exist, and the current has a fixed direction. When $\Delta$ is much greater than $J$, the current disappears in the bulk region, where the Néel order is maintained.
@@ 5.2 Effect of interaction in the Domain-Wall melting for Hatano-Nelson
Figure 11: Magnetisation profile in function of the lattice site and time. Left - Dynamics in Model A. Right - Dynamics in Model B. The system was initially prepared in a Néel State (Eq. (33)). Other parameters: $\Delta=0$, $\gamma=0.8J$ and $L=20$.
@@ 6 Application to Quantum Jumps Unravelling
The quantum jump is applied by first selecting the quantum jump channel $\mu$ in accordance to their probability
@@
This algorithm gives access to the full Lindbladian dynamics, by averaging over the propagated quantum trajectories. Further, it also provides access to the dynamics under continuous monitoring [20, 9]. This is accomplished by tracking the many-body quantum trajectory and computing, for example, non-linear functions of the state. One such function is the entanglement entropy of quantum trajectories, as we shall discuss below.
@@ 6.1 Magnetisation and Entanglement Dynamics in Monitored Spin Chains
where $J$ is the XX exchange term between neighbouring spins and $\Delta$ an Ising like exchange. We compare the dissipative dynamics generated by two different types of quantum jump operators. A first set of jump operators that we consider describe next-neighbour decoherence of spin excitations along the chain and take the form
@@
$L_{1+\ell}$ $=\sqrt{\left|\gamma\right|}\left(S^{-}_{\ell}-i\text{sgn}\left(\gamma\right)S^{-}_{\ell+1}\right),\;\ell=\{0,\cdots,L-2\},$
@@
Through the rest of the document, we name this system Model A. Interestingly, with this choice of jump operators the non-Hermitian Hamiltonian associated to the no-click limit turns out to be given by a spin version of the many-body Hatano-Nelson model. Indeed, the following straightforward calculation yields the non-Hermitian Hamiltonian
@@
The last term does not affect the dynamics, as it just an overall background decay. Although this model can be connected to the fermionic version of the Hatano-Nelson model, the Linblad equation is not quadratic, hence the quantum trajectories not gaussian, due to the Jordan-Wigner strings in the quantum jumps, i.e. terms of the form $S^{-}\rho S^{+}$.
@@
with =$\ell=\left\{0,\cdots,L-2\right\}$. We refer to this configuration as Model B. Previous studies have shown that these operators induce a phenomenon known as the Liouvillian skin effect [24, 108, 109, 110]. Specifically, there is an exponential localisation of the Liouvillian modes at the boundaries of the system. This becomes evident when the Linblad is projected onto the one-particle sector, where, at $J=0$ and $\Delta=0$, it simplifies to an effective Hatano-Nelson Hamiltonian tuned to its exceptional point [108],
@@
$\mathcal{H}_{\text{eff}}$ $=-\frac{J}{2}\sum_{\ell=0}^{L-2}\left[S_{\ell}^{+}S_{\ell+1}^{-}+\text{h.c}\right]-\left[\Delta-i\dfrac{\gamma}{2}\right]\sum_{\ell=0}^{L-2}S_{\ell}^{z}S_{\ell+1}^{z}$ (44)
$\quad+\frac{i\gamma}{4}\left(S_{0}^{z}-S_{L-1}^{z}\right)-\frac{i\gamma}{8}\left(L-1\right).$
@@
In Fig. 11, we plot the magnetisation dynamics starting from an initial Néel state and evolving under the two types of dissipative evolutions. In the left panel, we plot the dynamics generated by Model A (Eq. (40)). Under this set of jump operators, manifestations of non-reciprocity and charge accumulation are still present in the transient dynamics [111, 112]. We note that these are the same jump operators considered in [112] with the phase, $\phi$, defined by them, tuned to get the non-reciprocal regime. However, in the long time limit, a generic state converges to a configuration with zero excitations, $\ket{\downarrow\cdots\downarrow}$ [112] (see the left plot of Fig. 11). This is attributed to the recycling terms of the form $c_{n}\rho c^{\dagger}_{n}$, which effectively remove particles from the system. Additionally, the system has another dark state within the one-magnon sector, which only plays a role in decelerating the relaxation dynamics towards the fully polarised down state [112]. The final state obtained under this Linbladian does not resemble at all the magnetisation profile that one would have obtained in the no-click limit.
@@
We then focus on Model B (Eq. (42)), and plot in Fig. 11 (right panel) the resulting magnetisation dynamics. For the sake of simplicity, we consider the dynamics with $\Delta=0$. The dynamics driven by this model will facilitate the accumulation of spins in an up state at the left extremity of the chain. The imaginary Ising exchange term is responsible for diminishing the state’s norm when adjacent spins are antialigned, precipitating a quantum jump that propagates a spin towards the left edge. This phenomenon is depicted in the spatial and temporal evolution of the magnetisation profile shown in Fig. 11. In contrast to the behaviour observed in Model A, Model B does not converge to a steady state characterised by zero excitations. This distinction is attributable to a difference in symmetry. Model A exhibits only weak U(1) symmetry, whereas the Model B possesses strong U(1) symmetry, thereby ensuring that the system’s state is kept within the initial magnetisation sector at all times.
@@
Figure 12: Top - Time dependence of the conditional entanglement entropy for different system sizes with $\gamma=0.8J$. On the left, the dynamics corresponds to Model A, see Eq. (40)), while on the right to Model B, see Eq. (42). Bottom Left: Time evolution of the entanglement entropy for $\gamma=0.1J$ in the case of Model B. The inset corresponds to the fit of the steady-state entanglement entropy to the law $S_{L/2}\left(\infty\right)=a_{0}+a_{1}L$, with $a_{0}=\left(0.145\pm 0.002\right)$ and $a_{1}=\left(0.18\pm 0.03\right)$. Bottom Right - Steady-state entanglement entropy in Model B, for different system sizes as a function of the non-hermitian parameter. Other parameters: $\Delta=0.0J$.
@@
$\bar{S}_{\ell}\left(t\right)=\int\mathcal{D}\xi_{t}\;\mathcal{P}\left(\xi_{t}\right)S_{\ell}\left(\xi_{t}\right),$ (45)
@@
with $S\left(\xi_{t}\right)$ is the entanglement of a given quantum trajectory that evolves according to Eq. (1).
@@
For early times, the entanglement entropy in both models increases linearly. Without non-Hermiticity, the system evolves under standard hermitian dynamics, leading to the entanglement entropy saturating at a value proportional to the system’s volume as it locally thermalises [113, 114]. However, in the presence of non-Hermiticity and quantum jumps, this behaviour may drastically alter. In Model A, the entanglement entropy drops to zero after an initial period determined by the measurement rate $\gamma$ [115]. Each jump causes a spin-flip, driving the system to a state devoid of magnetic excitations. This occurs regardless of the system’s total size, resulting in a trivial entanglement area law for any nonzero value of $\gamma$. The no-click limit of Model A, which corresponds to the Hatano-Nelson model, also has this area law scaling of the entanglement entropy; however, this is driven by the single-particle skin effect[34].
@@
In contrast, the system in Model B does not relax to a zero excitation state, as seen in Fig. 12. The strong U(1) symmetry confines the dynamics to the magnetisation sector of the initial state [^3] The dynamics conserves the total number of excitations in the initial state as the initial state is eigenstate of $\sum_{n=0}^{L-1}S^{z}_{n}$ and the Hamiltonian in Eq. (44) as U(1) symmetry.. Initially, the system has a linear growth of the entanglement entropy, reminiscent of the unitary evolution. However, similarly to the Hatano-Nelson model, the steady-state supports an area law entanglement for a positive value of $\gamma$, as seen in the left plot of Fig.12. This can be understood through the spectral properties of the no-click Hamiltonian (Eq. (44)). The spectrum is always complex-valued, and so, in the no-click limit, the steady-state corresponds to right-eigenstate with the slowest decaying mode in the zero magnetisation sector. In particular, for certain values of $\gamma/J$, the imaginary component of the spectrum is gapped. Thus, the entanglement entropy inevitably follows an area law in the long-time limit, similar to the ground state of a gapped Hamiltonian [116, 32]. The imaginary gap ($\Delta_{\rm Im}$) in the zero magnetisation sector can be analytically obtained in the limit of $\gamma\gg J$,
$\Delta_{\rm Im}=-i\dfrac{\gamma}{2J}+\mathcal{O}\left(\dfrac{J}{\gamma}\right).$ (46)
@@
The measurement apparatus effectively disentangles the system, pushing excitations towards the left boundary, and preventing the formation of long-range correlations. It is harder to completely confirm this for smaller values of $\gamma$, where the unitary evolution dominates both the non-Hermitian and stochastic terms. This mainly affects systems with a smaller dimension, where finite-size effects are substantial, since the terms proportional to $L^{-n}$ with $n>0$ in the entanglement entropy cannot be ignored. For example, for $\gamma=0.1J$, one cannot extrapolate the true entanglement scaling, as the observed linear growth might be an artefact of small system sizes, so it remains uncertain whether a genuine entanglement transition occurs in the thermodynamic limit. Nevertheless, the remaining data points clearly indicate the collapse of all system sizes to the same value, thus revealing the area law nature of the entanglement entropy.